Check whether the matrices A and B are diagonalisable. Diagonalise those matrices which are diagonalisable. i) A =  1 0 0
1 2 −3
1 1 −2  
ii) B =  −2 −4 −1
3 5 1
1 1 2 .b) Find inverse of the matrix B in part a) of the question by using Cayley-Hamiltion theorem. c) Find the inverse of the matrix A in part a) of the question by finding its adjoint.
1
Expert's answer
2019-03-27T12:42:45-0400
"A=\begin{bmatrix}
 1 & 0 & 0 \\
 1 & 2 & -3 \\
1 & 1 & -2 \\
\end{bmatrix} , B=\begin{bmatrix}
 -2 & -4 & -1 \\
 3 & 5 & 1 \\
1 & 1 & 2 \\
\end{bmatrix}"

We first determine the eigenvalues of the matrix A


"|A|=1(2(-2)-(-3)(1))-0+0=-1""A-\lambda I=\begin{bmatrix}
 1-\lambda & 0 & 0 \\
 1 & 2-\lambda & -3 \\
1 & 1 & -2-\lambda \\
\end{bmatrix}"

Characteristic equation


"|A-\lambda I|=0""(1-\lambda)((2-\lambda)(-2-\lambda)-(-3)(1))-0+0=0""-(1-\lambda)^2(1+\lambda)=0""\lambda_1=1, \lambda_2=1, \lambda_3=-1."

These are eigenvalues.

Next, find the eigenvectors.


"\lambda=1""\begin{bmatrix}
 1-\lambda & 0 & 0 \\
 1 & 2-\lambda & -3 \\
1 & 1 & -2-\lambda \\
\end{bmatrix} =\begin{bmatrix}
 0 & 0 & 0 \\
 1 & 1 & -3 \\
1 & 1 & -3 \\
\end{bmatrix}"

Perform row operations to obtain the rref of the matrix:


"\begin{bmatrix}
 0 & 0 & 0 \\
 1 & 1 & -3 \\
1 & 1 & -3 \\
\end{bmatrix} \to \begin{bmatrix}
 1 & 1 & -3 \\
 0 & 0 & 0 \\
0 & 0 & 0 \\
\end{bmatrix}"

Now, solve the matrix equation


"\begin{bmatrix}
 1 & 1 & -3 \\
 0 & 0 & 0 \\
0 & 0 & 0 \\
\end{bmatrix} \begin{bmatrix}
 v_1 \\
 v_2 \\
v_3 \\
\end{bmatrix} =\begin{bmatrix}
 0 \\
 0 \\
0 \\
\end{bmatrix}"

If we take


"v_2=t, v_3=s, then\ v_1=3s-t."

Therefore


"v=\begin{bmatrix}
 3s-t \\
 t \\
3 \\
\end{bmatrix} =\begin{bmatrix}
 -1 \\
 1 \\
0 \\
\end{bmatrix}t +\begin{bmatrix}
 3 \\
 0 \\
1 \\
\end{bmatrix}s"

"\lambda=-1""\begin{bmatrix}
 1-\lambda & 0 & 0 \\
 1 & 2-\lambda & -3 \\
1 & 1 & -2-\lambda \\
\end{bmatrix} =\begin{bmatrix}
 2 & 0 & 0 \\
 1 & 3 & -3 \\
1 & 1 & -1 \\
\end{bmatrix}"

Perform row operations to obtain the rref of the matrix:


"\begin{bmatrix}
 2 & 0 & 0 \\
 1 & 3 & -3 \\
1 & 1 & -1 \\
\end{bmatrix} \to \begin{bmatrix}
 1 & 0 & 0 \\
 0 & 1 & -1 \\
0 & 0 & 0 \\
\end{bmatrix}"

Now, solve the matrix equation


"\begin{bmatrix}
 1 & 0 & 0 \\
 0 & 1 & -1 \\
0 & 0 & 0 \\
\end{bmatrix} \begin{bmatrix}
 v_1 \\
 v_2 \\
v_3 \\
\end{bmatrix} =\begin{bmatrix}
 0 \\
 0 \\
0 \\
\end{bmatrix}"

If we take


"v_3=t, then\ v_1=0, v_2=t."

Therefore


"v=\begin{bmatrix}
 0 \\
 t \\
t \\
\end{bmatrix} =\begin{bmatrix}
 0 \\
 1 \\
1 \\
\end{bmatrix}t"

Form the matrix P, whose i-th column is the i-th eigenvector:


"P=\begin{bmatrix}
 -1 & 3 & 0 \\
 1 & 0 & 1 \\
0 & 1 & 1 \\
\end{bmatrix}"

Form the diagonal matrix D, whose element at row i, column i is i-th eigenvalue:


"D=\begin{bmatrix}
 1 & 0 & 0 \\
 0 & 1 & 0 \\
0 & 0 & -1 \\
\end{bmatrix}"

These matrices have the property that 


"A=PDP^{-1}"

We first determine the eigenvalues of the matrix B


"|B|=-2(5(2)-(1)(1))+4(3(2)-(1)(1))-1(3(1)-(1)(5))=4""B-\lambda I=\begin{bmatrix}
 -2-\lambda & -4 & -1 \\
 3 & 5-\lambda & 1 \\
1 & 1 & 2-\lambda \\
\end{bmatrix}"

Characteristic equation


"|B-\lambda I|=0""(-2-\lambda)(5-\lambda)(2-\lambda)-3-4+(5-\lambda)-(-2-\lambda)+12(2-\lambda)=0""-(\lambda)^3+5(\lambda)^2-8\lambda+4=0""\lambda_1=1, \lambda_2=2, \lambda_3=2."

These are eigenvalues.

Next, find the eigenvectors.


"\lambda=1""\begin{bmatrix}
 -2-\lambda & -4 & -1 \\
 3 & 5-\lambda & 1 \\
1 & 1 & 2-\lambda \\
\end{bmatrix} =\begin{bmatrix}
 -3 & -4 & -1 \\
 3 & 4 & 1 \\
1 & 1 & 1 \\
\end{bmatrix}"

Perform row operations to obtain the rref of the matrix:


"\begin{bmatrix}
 -3 & -4 & -1 \\
 3 & 4 & 1 \\
1 & 1 & 1 \\
\end{bmatrix} \to \begin{bmatrix}
 1 & 0 & 3 \\
 0 & 1 & -2 \\
0 & 0 & 0 \\
\end{bmatrix}"

Now, solve the matrix equation


"\begin{bmatrix}
 1 & 0 & 3 \\
 0 & 1 & -2 \\
0 & 0 & 0 \\
\end{bmatrix} \begin{bmatrix}
 v_1 \\
 v_2 \\
v_3 \\
\end{bmatrix} =\begin{bmatrix}
 0 \\
 0 \\
0 \\
\end{bmatrix}"

If we take


"v_3=t, then\ v_1=-3t, v_2=2t."

Therefore


"v=\begin{bmatrix}
 -3t \\
 2t \\
t \\
\end{bmatrix} =\begin{bmatrix}
 -3 \\
 2 \\
1 \\
\end{bmatrix}t"

"\lambda=2""\begin{bmatrix}
 -2-\lambda & -4 & -1 \\
 3 & 5-\lambda & 1 \\
1 & 1 & 2-\lambda \\
\end{bmatrix} =\begin{bmatrix}
 -4 & -4 & -1 \\
 3 & 3 & 1 \\
1 & 1 & 0 \\
\end{bmatrix}"

Perform row operations to obtain the rref of the matrix:


"\begin{bmatrix}
 -4 & -4 & -1 \\
 3 & 3 & 1 \\
1 & 1 & 0 \\
\end{bmatrix} \to \begin{bmatrix}
 1 & 1 & 0 \\
 0 & 0 & 1 \\
0 & 0 & 0 \\
\end{bmatrix}"

Now, solve the matrix equation


"\begin{bmatrix}
 1 & 1 & 0 \\
 0 & 0 & 1 \\
0 & 0 & 0 \\
\end{bmatrix} \begin{bmatrix}
 v_1 \\
 v_2 \\
v_3 \\
\end{bmatrix} =\begin{bmatrix}
 0 \\
 0 \\
0 \\
\end{bmatrix}"

If we take


"v_2=t, then\ v_1=-t, v_3=0."

Therefore


"v=\begin{bmatrix}
 -t \\
 t \\
0 \\
\end{bmatrix} =\begin{bmatrix}
 -1 \\
 1 \\
0 \\
\end{bmatrix}t"

Since the number of eigenvectors is less than dimension of the matrix, then the matrix is not diagonalizable.

The matrix B is not diagonalizable.


b) Find inverse of the matrix B in part a) of the question by using Cayley-Hamiltion theorem.


"B=\begin{bmatrix}
 -2 & -4 & -1 \\
 3 & 5 & 1 \\
1 & 1 & 2 \\
\end{bmatrix}"

"|B|=-2(5(2)-(1)(1))+4(3(2)-(1)(1))-1(3(1)-(1)(5))=4\mathrlap{\,/}{=}0"


We have


"B-\lambda I=\begin{bmatrix}
 -2-\lambda & -4 & -1 \\
 3 & 5-\lambda & 1 \\
1 & 1 & 2-\lambda \\
\end{bmatrix}"

"p(\lambda)=|B-\lambda I|="

"=(-2-\lambda)(5-\lambda)(2-\lambda)-3-4+(5-\lambda)-(-2-\lambda)+12(2-\lambda)="

"=-(\lambda)^3+5(\lambda)^2-8\lambda+4"

Thus, we have obtained the characteristic polynomial


"p(\lambda)=-(\lambda)^3+5(\lambda)^2-8\lambda+4"

of the matrix B.

The Cayley-Hamilton theorem yields that


"0=p(B)=-B^3+5B^2-8B+4I"

where O is the 3×3 zero matrix.

Rearranging terms, we have


"B^3-5B^2+8B=4I""B({1 \over 4}(B^2-5B+8I))=I"

Similarly, we have


"({1 \over 4}(B^2-5B+8I))B=I"

It follows from these two equalities that the matrix


"{1 \over 4}(B^2-5B+8I)"

is the inverse matrix of B.

Therefore, we have


"B^{-1}={1 \over 4}(B^2-5B+8I)"

"B^2=\begin{bmatrix}
 -2 & -4 & -1 \\
 3 & 5 & 1 \\
1 & 1 & 2 \\
\end{bmatrix}\begin{bmatrix}
 -2 & -4 & -1 \\
 3 & 5 & 1 \\
1 & 1 & 2 \\
\end{bmatrix}="

"=\begin{bmatrix}
 -9 & -13 & -4 \\
 10 & 4 & 4 \\
3 & 3 & 4 \\
\end{bmatrix}"

"B^2-5B+8I=\begin{bmatrix}
 -9 & -13 & -4 \\
 10 & 4 & 4 \\
3 & 3 & 4 \\
\end{bmatrix}-5\begin{bmatrix}
 -2 & -4 & -1 \\
 3 & 5 & 1 \\
1 & 1 & 2 \\
\end{bmatrix}+"

"+8\begin{bmatrix}
 1 & 0 & 0 \\
 0 & 1 & 0 \\
0 & 0 & 1 \\
\end{bmatrix}=\begin{bmatrix}
 9 & 7 & 1 \\
 -5 & -3 & -1 \\
-2 & -2 & 2 \\
\end{bmatrix}"

Then the inverse matrix of B is


"B^{-1}=\begin{bmatrix}
 9/4 & 7/4 & 1/4 \\
 -5/4 & -3/4 & -1/4 \\
 -1/2 & -1/2 & 1/2 \\
\end{bmatrix}"



c) Find the inverse of the matrix A in part a) of the question by finding its adjoint.


"B=\begin{bmatrix}
 -2 & -4 & -1 \\
 3 & 5 & 1 \\
1 & 1 & 2 \\
\end{bmatrix}""|B|=-2(5(2)-(1)(1))+4(3(2)-(1)(1))-1(3(1)-(1)(5))=4\mathrlap{\,/}{=}0"


"C_{11}=\begin{vmatrix}
 5 & 1 \\
 1 & 2
\end{vmatrix}=9, C_{12}=-\begin{vmatrix}
 3 & 1 \\
 1 & 2
\end{vmatrix}=-5, C_{13}=\begin{vmatrix}
 3 & 5 \\
 1 & 1
\end{vmatrix}=-2,"




"C_{21}=-\begin{vmatrix}
 -4 & -1 \\
 1 & 2
\end{vmatrix}=7, C_{22}=\begin{vmatrix}
 -2 & -1 \\
 1 & 2
\end{vmatrix}=-3, C_{23}=-\begin{vmatrix}
 -2 & -4 \\
 1 & 1
\end{vmatrix}=-2,"




"C_{31}=\begin{vmatrix}
 -4 & -1 \\
 5 & 1
\end{vmatrix}=1, C_{32}=-\begin{vmatrix}
 -2 & -1 \\
 3 & 1
\end{vmatrix}=-1, C_{33}=\begin{vmatrix}
 -2 & -4 \\
 3 & 5
\end{vmatrix}=2."

"Adj(B)=C^T=\begin{bmatrix}
 9 & 7 & 1 \\
 -5 & -3 & -1 \\
-2 & -2 & 2 \\
\end{bmatrix}"

"B^{-1}={1 \over {|B|}}C^T=\begin{bmatrix}
 9/4 & 7/4 & 1/4 \\
 -5/4 & -3/4 & -1/4 \\
 -1/2 & -1/2 & 1/2 \\
\end{bmatrix}"


"B^{-1}=\begin{bmatrix}
 9/4 & 7/4 & 1/4 \\
 -5/4 & -3/4 & -1/4 \\
 -1/2 & -1/2 & 1/2 \\
\end{bmatrix}"